• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name spiegel/Techs

    Log entry time 12:23:20 on April 10, 2008

    Entry number 231178

    keyword=Move LEFT Spectrometer

    Moved spectrometer to 12.5 degrees. Found some interferance with lumi detector wiring and exit beam pipe backing pump tubing. Relocated wires and vacuum tube to make clearance for small angle at beam dump pipe support. Thank you for your help John Segal. That is all.

    A copy of this log entry has been emailed to: folts@jlab.org,segal@jlab.org